OpenStudy (anonymous):

Please help? An equation is shown below: 4(x - 3) - 5(x + 1) = 3 Which statement shows a correct next step in solving the equation? The equation can become 4x - 3 - 5x + 1 = 3 by applying the associative property of multiplication. The equation can become 4x - 3 - 5x + 1 = 3 by applying the distributive property. The equation can become 4x - 12 - 5x - 5 = 3 by applying the distributive property. The equation can become 4x - 12 - 5x - 5 = 3 by applying the associative property of multiplication.

OpenStudy (anonymous):

the last one

OpenStudy (ivettef365):

what property are you applying ?

OpenStudy (anonymous):

The is the last statement.

OpenStudy (anonymous):

answer*

OpenStudy (anonymous):

We have : 4x-12-5x-5=3 -x -23=3 Now answer !:) :D

OpenStudy (ivettef365):

@paul1231 you have the answer on the equation correct, 4x-12-5x-5=3, but which property did you use ??

OpenStudy (anonymous):

associative?

OpenStudy (anonymous):

or is it distributive, i'm really not sure

OpenStudy (ivettef365):

a+(b+c) = (a+b)+c this is associative a(b+c) = ab+ac this is distributive

OpenStudy (ivettef365):

or a.(b.c) = (a.b).c associative does your problem looks like that or like a(b+c) you have 4(x - 3)